I have a question that states: for the pair of functions, evaluate (g of f)(5). f(x)=x-9/x, g(x)=x^2+3. How do I do this?
Answer by stanbon(48516)

evaluate (g of f)(5). f(x)=(x-9)/x, g(x)=x^2+3.
--------------
g[f(5)] = g[(5-9)/5)]
=g[-4/5]
= (4/5)^2 + 3
= 16/25 = 75/25
=91/25
==============
Cheers,
stan H.